A Company sells a particular model of component, with most of the sales being made in the fall season. Company makes a one-time purchase of the components prior to each fall season at a cost of $40 for each component and sells for $60 each of them. Any component unsold at the end of the fall season are marked down to $29 and sold in a special winter sale. Sales of the past 10 season are given belo1. A Company sells a particular model of component, with most of the sales being made in the fall season. Company makes a onew:
Answer #1

ANSWER:

a)

Assuming demand follows a normal distribution, we can calculate the mean (m) and standard deviation (s) of the above data set and use that to arrive at an optimal order value based on the target service levels.

Year 1 2 3 4 5 6 7 8 9 10
Sales 30 50 30 60 10 40 30 30 20 40
Mean 34 Std Dev 14

The mean of the above set is 34 with a standard deviation of approximately 14. For service level calculation we need to find the cost of understocking = Unit Profit = Cs = 60 - 40 = 20. Cost of Over stocking is the purchase price net of salvage value = Co = 40 - 29 = 11. The service level for optimal profit is then given by Cs/(Cs+Co) = 20/31 = 64%. The z value corresponding to this service level is approximately z= 0.37. Hence the order quantity should be m +s*z = 39

b)

Since the number of units are finite over the last 10 years we can approximate it to the distribution of 10 to 60 units with the frequencies from the given table and apply the same principle of service levels as calculated above. The frequency table has been reconstructed below

Demand Frequency Probability Cumulative Probability
10 1 0.10 0.10
20 1 0.10 0.20
30 4 0.40 0.60
40 2 0.20 0.80
50 1 0.10 0.90
60 1 0.10 1.00
Total 10

Here the optimal service level at 64% corresponds to the demand at 40 (0.6 < 0.64 < 0.8). Hence the order quantity calculated by this method is about 40 which we see is close to the value calculated by approximating a normal distribution.

c)

The approximation of a normal distribution provides a very close value to the second method so is probably a good approximation. Also we observe the trend in the distribution is also centrally oriented with 30 and 40 being the central values.

NOTE:: I HOPE YOUR HAPPY WITH MY ANSWER....***PLEASE SUPPORT ME WITH YOUR RATING...

***PLEASE GIVE ME "LIKE"...ITS VERY IMPORTANT FOR ME NOW....PLEASE SUPPORT ME ....THANK YOU